Sue and Tessa's ages add up to 88 years an Sue is four yeas older than Tessa. How old is Sue?

Answer by ikleyn(52810)

.
Let Sue be x years old; then Tessa is (x-4) years old, according to the condition.


x + (x - 4) = 88


2x - 4 = 88

2x = 88 + 4

2x = 92

x = 92/2 = 46.      ANSWER